OpenStudy (anonymous):

HEEEELLLLLPPPPPP!!!!!!!!!!!!!!!!! there is the same number of men and women on the city bus when it departs from the station.At Grubb Street,8 men get off the bus and nobody gets on.After the Grubb Street stop,there are twice as many women as men on the bus.How many women are on the bus? A.4 B.6 C.8 D.12 E.16

OpenStudy (anonymous):

CAN SOMEONE HELP ME AND SHOW ME HOW TO DO THIS

OpenStudy (anonymous):

lets do this

hero (hero):

Okay, well I already know what the answer is. For you, you'll have to figure it out by trial and error or through algebra. I tried setting this up with algebra but I ran into an issue, but anyway, What you do is start with the answer you believe is correct. In this case, I propose that 16 is the answer. If 16 is the answer then there were 16 men and 16 women on the bus to begin with. Then 8 men got off the bus leaving 16 women and 8 men.

hero (hero):

I messed up this explanation really bad but there ya go.

OpenStudy (anonymous):

hmmm

hero (hero):

Do you get it now or are you still lost?

OpenStudy (anonymous):

im totally lost

OpenStudy (anonymous):

did you added 8+8 since they said twice as many

hero (hero):

Before Grubb Street Bus Stop: 16 women and 16 men After Grubb Street Bus Stop: (after 8 men get off) 16 women and 8 men There are twice as many women because 8 x 2 = 16

OpenStudy (anonymous):

but hoe did you got 16

OpenStudy (anonymous):

at first

hero (hero):

From the answer choices. I started with 16 because it's the only one that is a "double"

hero (hero):

other than 4, but 4 is too small

OpenStudy (anonymous):

ohhhhhh i get you

hero (hero):

actually 4, 8, and 16 are doubles.

OpenStudy (anonymous):

6 too

hero (hero):

But if you take away 8 you end up with 0 and you can't multiply 0 times anything and get "double"

hero (hero):

Actually, all even numbers are doubles

hero (hero):

2 + 2 = 4 3 + 3 = 6 4 + 4 = 8 6 + 6 = 12 8 + 8 = 16

hero (hero):

But 16 is the only one that works for this problem.

OpenStudy (anonymous):

So lets say we picked 8 What would we get or do

OpenStudy (anonymous):

there would be 8 women and 8 men then 8 men gets off but you cant do that becuase there will be no men am i correct

OpenStudy (anonymous):

then 8x0=0 then that wouldnt work

hero (hero):

I think you're getting the hang of it

OpenStudy (anonymous):

im LOVING THIS....THANK YOU im sooo happy

OpenStudy (anonymous):

it took me hours to figure this out

hero (hero):

It shouldn't have :P

OpenStudy (anonymous):

i know its that im getting ready for my SAT an these questions are soo hard,well some of them

hero (hero):

That was an SAT question? Too easy.

hero (hero):

I'm pretty sure the real ones will be more difficult.

OpenStudy (anonymous):

Dude no its not

hero (hero):

If it takes you hours to figure out that easy question, you should be worried.

OpenStudy (anonymous):

hahahahahaha dude your making worried.

OpenStudy (anonymous):

me

OpenStudy (anonymous):

dude im terrible at math and your like a genious am i right

hero (hero):

Yeah, because the SAT questions will be much more difficult

hero (hero):

I wouldn't say genius. I'm just really good at figuring things out. You have to at least know up to pre-calculus level to take SAT Math.

OpenStudy (anonymous):

how old are you BTW

hero (hero):

Why do you want to know that? :P

OpenStudy (anonymous):

dude just wondering but you dont have to answer my question its cool

OpenStudy (anonymous):

like if you went college did you did good on your SAT. Bet you did